OpenStudy (anonymous):

Suppose that water is pouring into a swimming pool in the shape of a right circular cylinder at a constant rate of 5 cubic feet per minute. If the pool has radius 7 feet and height 9 feet, what is the rate of change of the height of the water in the pool when the depth of the water in the pool is 7 feet?

OpenStudy (anonymous):

this is what i tried. dint work

OpenStudy (anonymous):

I can suggest something if you want.

OpenStudy (anonymous):

sure!

OpenStudy (anonymous):

Something between your work and question don't match up.....

OpenStudy (anonymous):

Is this supposed to be 2 questions?

OpenStudy (anonymous):

oh shoot

OpenStudy (anonymous):

never mind thanta for a totally different question

OpenStudy (anonymous):

thnx for pointing out

OpenStudy (anonymous):

But, do you need help for the first question?

OpenStudy (anonymous):

yes

OpenStudy (anonymous):

V is a funtion of t V(t) = 5t now think of h as a funtion of t h(t) h(t) = ((pi) r^2)/V(t) => h(t) = (pir^2)/5t set h(t) = 7, find t differentiate h(t) and plug in t

OpenStudy (anonymous):

V = pi * r^2 * h

OpenStudy (anonymous):

V = pi * 7^2 * h

OpenStudy (anonymous):

V = pi * 49h

OpenStudy (anonymous):

dV/dh = 49pi

OpenStudy (anonymous):

dV/dt * dt/dh = 49pi

OpenStudy (anonymous):

5 * dt/dh = 49pi

OpenStudy (anonymous):

dt/dh = 49pi/5

OpenStudy (anonymous):

dh/dt = 5/(49pi) =

OpenStudy (anonymous):

I might have made a mistake somewhere, correct me if I did.

OpenStudy (anonymous):

=.32057 ill check

OpenStudy (anonymous):

will the units be ft/min?

OpenStudy (anonymous):

yes

OpenStudy (anonymous):

its incorrect

OpenStudy (anonymous):

Ya, I just noticed I made a mistake

OpenStudy (anonymous):

one sec.

OpenStudy (anonymous):

@zordoloom did you realize your mistake?

OpenStudy (anonymous):

Ya, The h wasn't calculated properly.

OpenStudy (anonymous):

you mean 7^2 pi?

OpenStudy (anonymous):

That part, yes

OpenStudy (anonymous):

@zordoloom do you know what it should be?

OpenStudy (anonymous):

=.017897 ill check if its right

OpenStudy (anonymous):

incorrect

OpenStudy (anonymous):

h' was supposed to be the final answer right?

OpenStudy (anonymous):

wait

OpenStudy (anonymous):

it;s (5*27)/(49^2 pi) right?

OpenStudy (anonymous):

i'm taking a screen shot. u tell me if i put it in correct. wait

OpenStudy (anonymous):

did you get it?

OpenStudy (anonymous):

i'll try that wait

OpenStudy (anonymous):

did u get it? its incorrect

OpenStudy (anonymous):

i guess. thankyou very much for trying

OpenStudy (anonymous):

would the units be feet/min for sure? maybe thats whats wrong

OpenStudy (anonymous):

yeah i know. rounding to decimals for sure

OpenStudy (anonymous):

V=Ah then dh/dt=dv/dt/A dh/dt=5/7^2pi

OpenStudy (anonymous):

let me try it

OpenStudy (anonymous):

=.3205706789 did not work

OpenStudy (anonymous):

0.032057

OpenStudy (anonymous):

not correct

OpenStudy (anonymous):

try 5/(7^2pi)ft/min. i am sure this is correct.

OpenStudy (anonymous):

YES YES YES YES YES!!!!!

OpenStudy (anonymous):

lifesaver @wynnwynn

OpenStudy (anonymous):

Thank you very very much

OpenStudy (anonymous):

:)